DETECTION OF HI-Ar BY (2+1) REMPI SPECTROSCOPY

T. SUZUKI, H. KATAGANAGI, Institute for Molecular Science, Myodaiji, Okazaki 444, Japan; \underline{M. C. HEAVEN}, Department of Chemistry, Emory University, Atlanta, GA 30322.

ArHI was detected by spectroscopic means for the first time. A vibrationally resolved (2+1) REMPI spectrum was recorded for the g^{3}Σ^{−}(0^{+}) - X^{1}Σ^{+}(0^{+}) transition. Analysis of this data yielded stretching vibrational constants for the g state, and estimates for the van der Waals bond strengths in the g and X states. The ground state is weakly bound, and ab initio calculations predict a linear HI-Ar geometry. The implications of these results for studies of the one-atom cage effect in RgHX complexes are considered.